# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 5.3. Detection of embedded Klein bottles
First, we suppose that $M$ satisfies condition (i) in Lemma 5.6, i.e. $M$ is a closed $Sol$-manifold which is not fibered over $S^{1}$. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, it follows from [40, Corollary 4.12] that $N$ is also closed, and [54, Theorem 8.4] shows that $N$ also admits $Sol$-geometry. In addition, [31, Corollary 1.2] shows that the profinite completion of fundamental group of a compact, orientable 3-manifold determines whether this 3-manifold fibers over $S^{1}$. Thus, $N$ is also non-fibered, and satisfies condition (i) of Lemma 5.6.
Next, we suppose that $M$ satisfies condition (ii) in Lemma 5.6. Similarly, [54, Theorem 8.4] implies that $N$ is not a $Sol$-manifold. [56, Theorem B] shows that the profinite completion of fundamental group detects the JSJ-decomposition of a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary, which does not admit $Sol$-geometry. In fact, [56, Theorem B] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see for example [57, Theorem 5.6]. In particular, there is a one-to-one correspondence between the Seifert fibered pieces of $M$ and $N$, such that the corresponding pieces are profinitely isomorphic. Since one of the Seifert pieces of $M$ has a base orbifold which is either non-orientable or contains two singular points of index $2$, and this Seifert piece is obviously boundary incompressible and aspherical, it follows from Lemma 5.7 that the corresponding Seifert piece of $N$ also has a base orbifold which is either non-orientable or contains two singular points of index $2$. Therefore, $N$ also satisfies condition (ii) in Lemma 5.6. ∎
@@
The aspherical condition is important. In fact, there exist two lens spaces with isomorphic fundamental groups, such that one of them contains a Klein bottle while the other one does not; see [7, Corollary 6.4].
@@ 5.4. Profinite detection of exceptional Dehn fillings
Suppose $M$ and $N$ are orientable finite-volume hyperbolic 3-manifolds with cusps, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Let $\partial_{i_{0}}M$ be a boundary component of $M$. Then there exists a boundary component $\partial_{j_{0}}N$ of $N$, and an isomorphism $\psi:\pi_{1}\partial_{i_{0}}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j_{0}}N$, such that
@@
According to Proposition 2.2, $M_{c}$ is finite-volume hyperbolic if and only if $N_{\overline{\psi}(c)}$ is finite-volume hyperbolic. Hence, by definition of exceptional Dehn filling, $c\in\mathcal{E}(M,\partial_{i_{0}}M)$ if and only if $\overline{\psi}(c)\in\mathcal{E}(N,\partial_{j_{0}}N)$, which proves (1).
Since $M_{c}$ and $N_{\overline{\psi}(c)}$ have the same number of boundary components, Corollary 5.3 implies that $M_{c}$ is spherical if and only if $N_{\overline{\psi}(c)}$ is aspherical, which proves (2).
Similarly, (3) follows from Lemma 5.4 since the isomorphism $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ respects the peripheral structure, and (4) follows from Lemma 5.8. ∎
@@ 6.1. Some Dehn-surgeries on the Whitehead link
We start from the example of $\mathcal{W}(-2)$. Let $N$ be a compact orientable 3-manifold so that $\widehat{\pi_{1}\mathcal{W}(-2)}\cong\widehat{\pi_{1}N}$. Then Proposition 2.2 implies that $N$ is also finite-volume hyperbolic with exactly one cusp. Therefore, Proposition 5.10 implies that there is an isomorphism $\psi:\pi_{1}(\partial\mathcal{W}(-2))\to\pi_{1}(\partial N)$, such that $\overline{\psi}(\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2)))=\mathcal{E}_{t}(N,\partial N)$. Note that $\overline{\psi}$ is induced by a group isomorphism, so Lemma 6.2 implies that $\overline{\psi}$ preserves the geometric intersection number $\Delta(\cdot,\cdot)$. Thus, $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=\max\{\Delta(c_{1}^{\prime},c_{2}^{\prime})\mid c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2))\}=6$. According to Proposition 6.3 (1), the only possibility is $N\cong\mathcal{W}(-2)$.
@@ 6.2. Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link, and $M_{14}$
Recall that $\mathcal{W}$, $\mathcal{L}$ and $\mathcal{WS}$ denote the Whitehead link, the two-bridge link associated to the rational number $\frac{3}{10}$, and the Whitehead sister link (i.e. the $(-2,3,8)$-Pretzel link) respectively. $M_{14}$ is the double branched cover along the tangle $\mathcal{Q}$ in $S^{2}\times I$ (Figure 6) defined in [24].
Proposition 6.6 ([24, Corollary 1.3] ).
@@
For simplicity, we denote $M_{1}=S^{3}\setminus\mathring{N}(\mathcal{W})$, $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$, and $M_{3}=S^{3}\setminus\mathring{N}(\mathcal{WS})$ following the notation of [24]. It is well-known that $M_{1}$, $M_{2}$ and $M_{3}$ are finite-volume hyperbolic 3-manifolds, and [24, Lemma 23.13] proved that $M_{14}$ is also finite-volume hyperbolic.
@@
We first show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$. In fact, $M_{1}$ and $M_{2}$ are both hyperbolic link complements in $S^{3}$, while in the proof of [24, Lemma 24.2], Gordon-Wu showed that $M_{14}$ is not a link complement. Therefore, Theorem 4.1 implies $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$.
Next, we show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{2}}$. Suppose on the contrary that $f:\widehat{\pi_{1}M_{1}}\to\widehat{\pi_{1}M_{2}}$ is an isomorphism. Fix the conjugacy representatives of peripheral subgroups $\pi_{1}\partial_{i}M_{1}\subseteq\pi_{1}M_{1}$ and $\pi_{1}\partial_{i}M_{2}\subseteq\pi_{1}M_{2}$ ($i=1,2$). Then Theorem 3.2 (1) implies that there exists $i\in\{1,2\}$ such that $f(\overline{\pi_{1}\partial_{1}M_{1}})$ is a conjugate of $\overline{\pi_{1}\partial_{i}M_{2}}$ in $\widehat{\pi_{1}M_{2}}$. By composing with a conjugation, we may simply assume $f(\overline{\pi_{1}\partial_{1}M_{1}})=\overline{\pi_{1}\partial_{i}M_{2}}$. Recall that $\overline{\pi_{1}\partial_{i}M_{j}}\cong\widehat{\pi_{1}\partial_{i}M_{j}}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M_{j}$ by Proposition 2.3. Then the abelianization of profinite groups [57, Proposition 2.11] yields the following commutative diagram.
[…]
In fact, the vertical maps in the commutative diagram are induced by the inclusions $\rho:\partial_{1}M_{1}\to M_{1}$ and $\iota:\partial_{i}M_{2}\to M_{2}$, i.e. ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})\xrightarrow{1\otimes\rho_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{1};\mathbb{Z})$ and likewise.
Note that the linking number of the two components of $\mathcal{W}$ is 00, so each boundary component of $M_{1}=S^{3}\setminus\mathring{N}(\mathcal{W})$ contains a null-homologous boundary slope. In other words, $\rho_{\ast}:H_{1}(\partial_{1}M_{1};\mathbb{Z})\to H_{1}(M_{1};\mathbb{Z})$ is not injective. Note that $H_{1}(\partial_{1}M_{1};\mathbb{Z})$ is a finitely generated Abelian group, so $H_{1}(\partial_{1}M_{1};\mathbb{Z})$ injects into ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})$, and ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})\xrightarrow{1\otimes\rho_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{1};\mathbb{Z})$ is not injective. On the other hand, the linking number of the two components of $\mathcal{L}$ is ±$\pm 3$ (depending on the orientation). Thus, the two boundary components of $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$ are both $H_{1}$-injective, i.e. $\iota_{\ast}:H_{1}(\partial_{i}M_{2};\mathbb{Z})\to H_{1}(M_{2};\mathbb{Z})$ is injective. Since $\widehat{\mathbb{Z}}$ is torsion-free, and hence a flat $\mathbb{Z}$-module, the homomorphism ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{i}M_{2};\mathbb{Z})\xrightarrow{1\otimes\iota_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{2};\mathbb{Z})$ is injective. Therefore, the following diagram shall not commute, which yields a contradiction.
[…]
@@ 6.3. More hyperbolic knot complements in $S^{3}$
    The Eudave-Muñoz knots $K(3,1,n,0)$ defined in [18] (Figure 4(a)) can be viewed as a $-\frac{1}{n}$-surgery on the unknotted component of the Whitehead sister link $\mathcal{WS}$ shown in Figure 1(b). Indeed, [18, Proposition 2.2] shows that $K(3,1,n,0)$ is hyperbolic for all $n\in\mathbb{Z}\setminus\{0\}$, and the mirror-image of $K(3,1,n,0)$ is $K(2,-1,1-n,0)$ by [18, Proposition 1.4].
Proposition 6.9 ([24, Theorem 24.4] ).
Suppose ${K}$ is a hyperbolic knot in $S^{3}$.
@@
    If there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$, then $K$ is the Eudave-Muñoz knot $K(3,1,n,0)$ or its mirror-image $K(2,-1,1-n,0)$ for some $n\in\mathbb{Z}\setminus\{0\}$.
@@
(1) and (2) are exactly (4) and (3) of [12, Theorem 1.1], and the calculation of geometric intersection number follows from Example 6.1. (3) is proven in [24, Corollary 24.5], see also [37].
Now we move on to (4). The fact that $\{25n-9,25n-\frac{13}{2}\}\subseteq\mathcal{E}_{t}(K(3,1,n,0))$ follows from [19, Proposition 5.4 (1) and (11)] [^3] In fact, there is a calculation error in the Seifert invariants in [19, Proposition 5.4 (1)], but the surgery coefficient is still correct., together with the fact that $K(3,1,n,0)$ is the mirror-image of $K(2,-1,1-n,0)$ [18, Proposition 1.4]. In particular, when $n=1$, $K(3,1,n,0)$ is the $(-2,3,7)$-Pretzel knot, and $\mathcal{E}_{t}(K(3,1,n,0))=\{16,\frac{37}{2},20\}$ strictly contains $\{25n-9,25n-\frac{13}{2}\}=\{16,\frac{37}{2}\}$. Nevertheless, any hyperbolic knot in $S^{3}$ has at most one non-integral toroidal slope as shown by [25, Theorem 1]. Thus, for any $n\in\mathbb{Z}\setminus\{0\}$, $\mathcal{E}_{t}(K(3,1,n,0))$ contains exactly one half-integral element, which is $25n-\frac{13}{2}$ stated above. ∎
@@
    For $m,n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$ if and only if $m=\pm n$; in this case, $\mathcal{J}_{m}$ and $\mathcal{J}_{n}$ are either equivalent or the mirror-image of each other.
@@
Similarly, if $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$, then $\{-2-9m,2-9m\}=\mathcal{E}_{t}(K)=\pm\mathcal{E}_{t}(K^{\prime})=\pm\{-2-9n,2-9n\}$ by Proposition 6.10, which implies $m=\pm n$. In addition, $\mathcal{J}_{n}$ can be viewed as a $\frac{1}{n}$-surgery on the link $\mathcal{L}$, while $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus, the $\frac{1}{n}$-surgery on $\mathcal{L}$ yields the mirror-image of the $-\frac{1}{n}$-surgery on $\mathcal{L}$. This implies that $\mathcal{J}_{n}$ and $\mathcal{J}_{-n}$ are mirror-image of each other. In particular, $X_{\mathcal{J}_{n}}$ is homeomorphic to $X_{\mathcal{J}_{-n}}$, so $\widehat{\pi_{1}X_{\mathcal{J}_{n}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{-n}}}$. This finishes the proof of (2).
@@
(4) is difficult to detect from the toroidal slopes; instead, we can detect these knots through the profinite rigidity of Alexander polynomials [46]. In fact, Ueki [46, Theorem 1.1] shows that if two knot complements $X_{K}$ and $X_{K^{\prime}}$ are profinitely isomorphic, then their Alexander polynomials $\Delta_{K}(t)$ and $\Delta_{K^{\prime}}(t)$ coincide up to a multiplicative unit in $\mathbb{Z}[t^{\pm}]$. According to [42, Chapter 7.B, Exercise 7], the Alexander polynomial of the twist knot $\mathcal{K}_{m}$ is $\Delta_{\mathcal{K}_{m}}(t)\doteq mt^{2}+(1-2m)t+m$. It is easy to verify that for any $m\neq n\in\mathbb{Z}$, $\Delta_{\mathcal{K}_{m}}(t)$ and $\Delta_{\mathcal{K}_{n}}(t)$ are not equivalent up to multiplicative unit, and therefore, $\widehat{\pi_{1}X_{\mathcal{K}_{m}}}\not\cong\widehat{\pi_{1}X_{\mathcal{K}_{n}}}$. ∎
@@
For brevity of notation, let $K$ denote one of the above listed knots, i.e. $K$ belongs to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), ${\mathcal{J}_{n}}$ ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and ${K(3,1,-n,0)}$ ($n\in\mathbb{Z}\setminus\{0\}$), and we denote $M=X_{K}$.
@@
When $K=\mathcal{K}_{n}$ or $K=\mathcal{J}_{n}$ for some $n\in\mathbb{Z}\setminus\{-1,0,1\}$, Proposition 6.10 (2) and (3) imply that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=4$. As a result, there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$, and Proposition 6.9 (2) implies that $K^{\prime}$ is either equivalent or mirror image to $\mathcal{K}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{0,1\}$) or $\mathcal{J}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{-1,0,1\}$). In addition, $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. According to (1), (2), (4) of Lemma 6.11, these candidates can be profinitely distinguished from each other, which ensures that $K^{\prime}$ can only be equivalent or mirror image to $K$. Thus, $N\cong X_{K^{\prime}}\cong X_{K}=M$, finishing the proof. ∎

### v3

@@ 5.3. Detection of embedded Klein bottles
First, we suppose that $M$ satisfies condition (i) in Lemma 5.7, i.e. $M$ is a closed $Sol$-manifold which is not fibered over $S^{1}$. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, it follows from [40, Corollary 4.12] that $N$ is also closed, and [53, Theorem 8.4] shows that $N$ also admits $Sol$-geometry. In addition, [31, Corollary 1.2] shows that the profinite completion of fundamental group of a compact, orientable 3-manifold determines whether this 3-manifold fibers over $S^{1}$. Thus, $N$ is also non-fibered, and satisfies condition (i) of Lemma 5.7.
Next, we suppose that $M$ satisfies condition (ii) in Lemma 5.7. Similarly, [53, Theorem 8.4] implies that $N$ is not a closed $Sol$-manifold. [55, Theorem B] shows that the profinite completion of fundamental group detects the JSJ-decomposition of a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary which does not admit $Sol$-geometry. In fact, [55, Theorem B] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see for example [56, Theorem 5.5]. In particular, there is a one-to-one correspondence between the Seifert fibered JSJ-pieces of $M$ and $N$, such that the corresponding pieces are profinitely isomorphic. Since one of the Seifert pieces of $M$ has a base orbifold which is either non-orientable or contains two singular points of index $2$, and this Seifert piece is obviously boundary incompressible and aspherical, it follows from Lemma 5.8 that the corresponding Seifert piece of $N$ also has a base orbifold which is either non-orientable or contains two singular points of index $2$. Therefore, $N$ also satisfies condition (ii) in Lemma 5.7. ∎
@@
The aspherical condition is important. In fact, there exist two lens spaces with isomorphic fundamental groups, such that one of them contains a Klein bottle while the other one does not; see [8, Corollary 6.4].
@@ 5.4. Profinite detection of exceptional Dehn fillings
Suppose $M$ and $N$ are orientable cusped finite-volume hyperbolic 3-manifolds, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Let $\partial_{i_{0}}M$ be a boundary component of $M$. Then, there exists a boundary component $\partial_{j_{0}}N$ of $N$, and an isomorphism $\psi:\pi_{1}\partial_{i_{0}}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j_{0}}N$, such that
@@
According to Proposition 2.3, $M_{c}$ is finite-volume hyperbolic if and only if $N_{\overline{\psi}(c)}$ is finite-volume hyperbolic. Hence, by definition of exceptional Dehn filling, $c\in\mathcal{E}(M,\partial_{i_{0}}M)$ if and only if $\overline{\psi}(c)\in\mathcal{E}(N,\partial_{j_{0}}N)$, which proves (1). Since $M_{c}$ and $N_{\overline{\psi}(c)}$ have the same number of boundary components, Corollary 5.3 implies that $M_{c}$ is aspherical if and only if $N_{\overline{\psi}(c)}$ is aspherical, which proves (2). Similarly, (3) follows from Lemma 5.5 since the isomorphism $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ respects the peripheral structure, and (4) follows from Lemma 5.9. ∎
@@ 6.1. Some Dehn-surgeries on the Whitehead link
We start from the example of $\mathcal{W}(-2)$. Let $N$ be a compact orientable 3-manifold so that $\widehat{\pi_{1}\mathcal{W}(-2)}\cong\widehat{\pi_{1}N}$. Then, Proposition 2.3 implies that $N$ is also finite-volume hyperbolic with exactly one cusp. Therefore, Proposition 5.11 implies that there is an isomorphism $\psi:\pi_{1}(\partial\mathcal{W}(-2))\to\pi_{1}(\partial N)$, such that $\overline{\psi}(\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2)))=\mathcal{E}_{t}(N,\partial N)$. Note that $\overline{\psi}$ is induced by a group isomorphism, so Lemma 6.2 implies that $\overline{\psi}$ preserves the geometric intersection number $\Delta(\cdot,\cdot)$. Thus, $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=\max\{\Delta(c_{1}^{\prime},c_{2}^{\prime})\mid c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2))\}=6$. According to Proposition 6.3 (1), the only possibility is $N\cong\mathcal{W}(-2)$.
@@ 6.2. Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link, and $M_{14}$
Recall that $\mathcal{W}$, $\mathcal{L}$ and $\mathcal{WS}$ denote the Whitehead link, the two-bridge link defined by the rational number $\frac{3}{10}$ in the Schubert normal form, and the Whitehead sister link which is the (2,3,8)(-2,3,8)-Pretzel link, respectively. $M_{14}$ is the double branched cover along the tangle $\mathcal{Q}$ in $S^{2}\times I$ (Figure 6) defined in [25].
Proposition 6.6 ([25, Corollary 1.3]).
@@
For simplicity, we denote $M_{1}=S^{3}\setminus\mathring{N}(\mathcal{W})$, $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$, and $M_{3}=S^{3}\setminus\mathring{N}(\mathcal{WS})$ following the notation of [25]. It is well-known that $M_{1}$, $M_{2}$ and $M_{3}$ are finite-volume hyperbolic 3-manifolds, and [25, Lemma 23.13] proved that $M_{14}$ is also finite-volume hyperbolic.
@@
We first show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$. In fact, $M_{1}$ and $M_{2}$ are both hyperbolic link complements in $S^{3}$, while in the proof of [25, Lemma 24.2], Gordon-Wu showed that $M_{14}$ is not a link complement. Therefore, Theorem 4.1 implies $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$.
Next, we show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{2}}$. Suppose by contrary that $\widehat{\pi_{1}M_{1}}\cong\widehat{\pi_{1}M_{2}}$. Then, by Theorem A, there exist a boundary component $\partial_{\ast}M_{1}$, a boundary component $\partial_{\star}M_{2}$, and a bijection $\overline{\psi}:\text{slp}(\partial_{\ast}M_{1})\to\text{slp}(\partial_{\star}M_{2})$ such that for any slope $c\in\text{slp}(\partial_{\ast}M_{1})$, $\widehat{\pi_{1}(M_{1})_{c}}\cong\widehat{\pi_{1}(M_{2})_{\overline{\psi}(c)}}$. In particular, $b_{1}((M_{1})_{c})=b_{1}((M_{2})_{\overline{\psi}(c)})$ by Lemma 2.1.
The boundary component $\partial_{\ast}M_{1}$ is given by a component $K_{\ast}$ of the Whitehead link $\mathcal{W}$. Now we let $c\in\text{slp}(\partial_{\ast}M_{1})$ be the longitude of $K_{\ast}$. Since the linking number of the two components in $\mathcal{W}$ is 0, the slope $c$ is actually null homologous in $M_{1}$. Thus, $b_{1}((M_{1})_{c})=b_{1}(M_{1})=2$. However, the linking number of the two components in $\mathcal{L}$ is $\pm 3$ which is non-zero, so both boundary components of $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$ are $H_{1}$-injective. Consequently, $b_{1}((M_{2})_{c^{\prime}})=1$ for any $c^{\prime}\in\text{slp}(\partial_{\star}M_{2})$. In particular, $b_{1}((M_{1})_{c})\neq b_{1}((M_{2})_{\overline{\psi}(c)})$, which yields a contradiction. ∎
@@ 6.3. More hyperbolic knot complements in $S^{3}$
    The Eudave-Muñoz knots $K(3,1,n,0)$ defined in [19] (Figure 4(a)) can be viewed as a $\frac{1}{n}-$surgery on the unknotted component of the Whitehead sister link $\mathcal{WS}$ shown in Figure 1(b). Indeed, [19, Proposition 2.2] shows that $K(3,1,n,0)$ is hyperbolic for all $n\in\mathbb{Z}\setminus\{0\}$, and the mirror-image of $K(3,1,n,0)$ is K(2,1,1n,0)K(2,-1,1-n,0) by [19, Proposition 1.4].
Proposition 6.9 ([25, Theorem 24.4]).
Suppose K{K} is a hyperbolic knot in $S^{3}$.
@@
    If there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$, then $K$ is the Eudave-Muñoz knot $K(3,1,n,0)$ or its mirror-image K(2,1,1n,0)K(2,-1,1-n,0) for some $n\in\mathbb{Z}\setminus\{0\}$.
@@
(1) and (2) are exactly (4) and (3) of [12, Theorem 1.1], and the calculation of geometric intersection number follows from Example 6.1. (3) is proven in [25, Corollary 24.5], see also [37].
Now we move on to (4). The fact that $\{25n-9,25n-\frac{13}{2}\}\subseteq\mathcal{E}_{t}(K(3,1,n,0))$ follows from [20, Proposition 5.4 (1) and (11)] [^1] In fact, there is a calculation error in the Seifert invariants in [20, Proposition 5.4 (1)], but the surgery coefficient is still correct., together with the fact that $K(3,1,n,0)$ is the mirror-image of K(2,1,1n,0)K(2,-1,1-n,0) [19, Proposition 1.4]. In particular, when $n=1$, $K(3,1,n,0)$ is the (2,3,7)(-2,3,7)-Pretzel knot, and $\mathcal{E}_{t}(K(3,1,n,0))=\{16,\frac{37}{2},20\}$ strictly contains $\{25n-9,25n-\frac{13}{2}\}=\{16,\frac{37}{2}\}$. Nevertheless, any hyperbolic knot in $S^{3}$ has at most one non-integral toroidal slope as shown by [26, Theorem 1]. Thus, for any $n\in\mathbb{Z}\setminus\{0\}$, $\mathcal{E}_{t}(K(3,1,n,0))$ contains exactly one half-integral element, which is $25n-\frac{13}{2}$ stated above. ∎
@@
    For $m,n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$ if and only if $m=\pm n$; in this case, $\mathcal{J}_{m}$ and $\mathcal{J}_{n}$ are either equivalent or the mirror images of each other.
@@
Similarly, if $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$, then $\{-2-9m,2-9m\}=\mathcal{E}_{t}(\mathcal{J}_{m})=\pm\mathcal{E}_{t}(\mathcal{J}_{n})=\pm\{-2-9n,2-9n\}$ by Proposition 6.10, which implies $m=\pm n$. In addition, $\mathcal{J}_{n}$ can be viewed as a $\frac{1}{n}$-surgery on the link $\mathcal{L}$, while $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus, the $\frac{1}{n}$-surgery on $\mathcal{L}$ yields the mirror-image of the $\frac{1}{n}-$surgery on $\mathcal{L}$. This implies that $\mathcal{J}_{n}$ and $\mathcal{J}_{-n}$ are the mirror images of each other. In particular, $X_{\mathcal{J}_{n}}$ is homeomorphic to $X_{\mathcal{J}_{-n}}$, so $\widehat{\pi_{1}X_{\mathcal{J}_{n}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{-n}}}$. This finishes the proof of (2).
@@
(4) is difficult to detect from the toroidal slopes; instead, we can detect these knots through the profinite rigidity of Alexander polynomials. Ueki [46, Theorem 1.1] shows that if two knot complements $X_{K}$ and $X_{K^{\prime}}$ are profinitely isomorphic, then their Alexander polynomials $\Delta_{K}(t)$ and $\Delta_{K^{\prime}}(t)$ coincide up to a multiplicative unit in $\mathbb{Z}[t^{\pm}]$. According to [42, Chapter 7.B, Exercise 7], the Alexander polynomial of the twist knot $\mathcal{K}_{m}$ is $\Delta_{\mathcal{K}_{m}}(t)\doteq mt^{2}+(1-2m)t+m$. It is easy to verify that for any $m\neq n\in\mathbb{Z}$, $\Delta_{\mathcal{K}_{m}}(t)$ and $\Delta_{\mathcal{K}_{n}}(t)$ are not equivalent up to multiplicative unit, and therefore, $\widehat{\pi_{1}X_{\mathcal{K}_{m}}}\not\cong\widehat{\pi_{1}X_{\mathcal{K}_{n}}}$. ∎
@@
For brevity of notation, let $K$ denote one of the above listed knots, i.e. $K$ belongs to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), $\mathcal{J}_{n}$} ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and K(3,1,n,0){K(3,1,-n,0)} ($n\in\mathbb{Z}\setminus\{0\}$), and we denote $M=X_{K}$.
@@
When $K=\mathcal{K}_{n}$ or $K=\mathcal{J}_{n}$ for some $n\in\mathbb{Z}\setminus\{-1,0,1\}$, Proposition 6.10 (2) and (3) imply that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=4$. As a result, there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$, and Proposition 6.9 (2) implies that $K^{\prime}$ is either equivalent or mirror image to $\mathcal{K}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{0,1\}$) or $\mathcal{J}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{-1,0,1\}$). According to (1), (2), (4) of Lemma 6.11, these candidates can be profinitely distinguished from each other, so $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$ ensures that $K^{\prime}$ can only be equivalent or mirror image to $K$. Thus, $N\cong X_{K^{\prime}}\cong X_{K}=M$, finishing the proof. ∎